Fermi National Laboratory


Accelerator Update for September 12, 2005 - September 14, 2005

Monday, September 12
The day shift began with Operations monitoring store 4385 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations put the Recycler into shot setup at 9:29 AM to stash antiprotons.

The stash ended and Pbar resumed stacking at 9:41 AM. MCenter resumed taking beam.

The Recycler prepared for stashing at 10:58 AM. The stash ended and Pbar and MCenter resumed normal operations at 11:58 AM.

Pbar experts began a reverse proton study at 12:37 PM.

The Pbar study ended and stacking resumed at 7:59 PM.

At 10:17 PM, Operations reported having a Pbar bunch rotation problem, which was caused by the heaters for a Debuncher RF station (DRF1) not regulating properly.

Tuesday, September 13
The midnight shift began with Operations monitoring store 4385 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Tevatron (TeV) experts began an end-of-store study at 7 AM.

TeV studies ended at 9:02 AM and Operations terminated store 4385 at 9:04 AM.

Operations put the TeV into shot setup at 9:18 AM.

Operations established combination store 4386 at 11:42 AM with an initial luminosity of 96.4E30. Pbar and MCenter resumed normal operations.

At 3:31 PM, quadrupole magnet power supplies tripped off for the Neutron Therapy Facility (NTF). Operators found its water skid tripped off due to a high temperature indication. Operators reset the system, but the magnet interlock system wouldn't reset. A NTF expert investigated.

Operations reported at 4:06 PM that RF experts were working on a Main Injector (MI) RF station (MIRF16) and their work was causing two other station (MIRF17 & 18) to trip off.

Operations reported at 4:20 PM that work on a MI kicker in MI-52 halted stacking.

Experts completed their work on the MI-52 kicker at 11:23 PM. Pbar resumed stacking.

The Recycler began stashing antiprotons at 11:43 PM.

Wednesday, September 14
The midnight shift began with Operations monitoring store 4386 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

The Recycler stash ended at 12:51 AM. Pbar and MCenter resumed normal operations.

At 1:08 AM, a Recycler power supply (R:CE621), a Booster power supply (WAPS) and RF, MIRF5 & 6, the Tevatron Electron Lens (TEL), and other equipment tripped off at the same time. The trips may have been caused by a lightning strike, but operators couldn't find any indication. Recycler lost some antiprotons because of this glitch.

At 5:50 AM, another glitch tripped off Recycler and caused the loss of its stash. Many of the systems that tripped on the previous trip, tripped off again, plus a few other systems. The High Voltage Coordinator said that the trips were likely due to off-site ComEd glitches caused by the weather.

The Plans for Wednesday
The plans for today are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/14/2005   email Fermilab